OpenStudy (anonymous):

If W is a non real cube root of unity then \[(1+W)(1+w^2)(2+W^4)(1+W^8)(1+w^{10})(1+w^{32})\] =?

OpenStudy (anonymous):

@mukushla @ghazi

OpenStudy (anonymous):

\[w^4 = w \]

OpenStudy (anonymous):

w^8 = w^2 w^10 = w w^32 = w^2

OpenStudy (anonymous):

(1+W)^3 + (1+w^2)^3

OpenStudy (anonymous):

1+w + w^2 = 0 1+w = -w^2

OpenStudy (anonymous):

(-w^2)^3 + (-w)^3

OpenStudy (anonymous):

and...i am stuck from here

OpenStudy (anonymous):

\[(1+w)(1+w^2)(1+w^4)(1+w^8)(1+w^{10})(1+w^{32})\]?

OpenStudy (anonymous):

Find the value

OpenStudy (anonymous):

\[w^2=-1-w\]\[w^4=w\]\[w^8=w^2=-1-w\]\[w^{10}=w\]\[w^{32}=w^2=-1-w\]

OpenStudy (anonymous):

Yup..

OpenStudy (anonymous):

i got this.....(-w^2)^3 + (-w)^3???

OpenStudy (anonymous):

is that 2 there (2+w^4)?

OpenStudy (anonymous):

1+w = -w^2

OpenStudy (anonymous):

(1+W)^3 + (1+w^2)^3

OpenStudy (anonymous):

\[(1+w)^2(1+w^2)^3(2+w)=(1+2w+w^2)(1-1-w)^3(2+w)\\=w\times(-w)^3\times(2+w)=-w(2+w)=-2w-w^2=-2w+1+w=1-w\]

OpenStudy (anonymous):

but i think thats (1+W^4) and answer will be 1

OpenStudy (anonymous):

yes the answer should be 1....

OpenStudy (anonymous):

but hw did u get that

OpenStudy (anonymous):

i got -2 as answer

OpenStudy (anonymous):

so with my notations\[(1+w)(1+w^2)(1+w^4)(1+w^8)(1+w^{10})(1+w^{32})\\=(1+w)^3(1+w^2)^3\]right?

OpenStudy (anonymous):

Yup...

OpenStudy (anonymous):

\[(1+w)^3(1+w^2)^3=(1+w)(1+w)^2(1-1-w)^3\\=(1+w)(1+2w+w^2)(-w)^3=-w(1+w)=-w-w^2=1\]

OpenStudy (anonymous):

sorry u r offline and i gave u the complete solution...see u yahoo